A 100-m3 open tank contains water 3 m deep. A. What is the pressure on the water at the bottom of the tank, relative to atmospheric pressure?
Answer #1

You can calculate the hydrostatic pressure of the liquid in a tank as the force per area for the area of the bottom of the tank as given by pressure = force/area units. In this case, the force would be the weight the liquid exerts on the bottom of the tank due to gravity.

Let base area of the tank is A square meter.

Because depth of water is =3meter

Therefore volume of water (V) = 3×A cubic meter

Density of water (d) in SI unit = 1000

So, mass of water = V×d

= 3×A×1000

= 3000×A kg

So total force on the base F= mg

= 3000×A×10 N

= 30000A Newtons

Therefore pressure at the base

P = F/A = 30000 A/ A = 30000 N/ square meter

= 3.0×10⁴ N/m²